OpenStudy (anonymous):

Solve the equation by completing the square. 6x^2-x-5=0 So far i got x^2-1/6x-5/6=0 x^2-1/6x=5/6 Now i dont understand this step by add the square of half of -1/6 to both sides I dont understand what that means please help

OpenStudy (anonymous):

why?....

pooja195 (pooja195):

Cant we just factor it?

OpenStudy (anonymous):

okay

pooja195 (pooja195):

@Tyler1204 ?

OpenStudy (anonymous):

@pooja195 i have the steps and it says add the square of half of -1/6 to both sides but i dont understand how that answer is 1/144

pooja195 (pooja195):

hmm im not sure how to do it that way @Loser66 might know :3

OpenStudy (anonymous):

This is what it says by the way http://prntscr.com/7dnq4a

OpenStudy (loser66):

you are good until \(x^2 -(1/6)x -(5/6)=0\) Now, I pick the middle term to say, ok? to have a complete the square, you have formula \((x-a)^2 = x^2 -2ax+a^2\) right? your middle term is (1/6) x, right?

OpenStudy (loser66):

so, to find a, the term we need to have the complete square, I do \(2a x = \dfrac{1}{6}x\), hence \(2a=\dfrac{1}{6}\), so \(a =\dfrac{1}{12}\) so far so good?

OpenStudy (anonymous):

erm erm @Loser66 ok like what you saying sounds good lol but like i have to follow this way :S did you see the picture i posted ? http://prntscr.com/7dnq4a

OpenStudy (loser66):

Now, the original one \(x^2 -2 (\dfrac{1}{12}) x +\color{red}{(\dfrac{1}{12})^2} - \color{red}{(\dfrac{1}{12})^2}-\dfrac{5}{6}=0\)

OpenStudy (loser66):

@Tyler1204 it is not "sounds good". I explained you why 1/144 appears.

OpenStudy (anonymous):

OH OH erm ok let me re read then lol sorry

OpenStudy (loser66):

You killed my feeling.

OpenStudy (anonymous):

Grrr im still confused because for me it dos not say anything about that formula you said :/ (x-a)^2=x^2-2ax+a^2 Did you open the picture? because like i have to follow what it says :S

OpenStudy (anonymous):

@Loser66